# FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation

### v1

@@ A.1 Parameter-Efficient Fine-Tuning (PEFT)
Fine-tuning LLMs is resource-intensive due to their large parameter counts. Parameter-efficient fine-tuning (PEFT) methods mitigate this by updating only a small subset of parameters while keeping pre-trained weights frozen Li and Liang (2021); He et al. (2021); Wang et al. (2022). Several PEFT approaches have been proposed, Adapter Tuning Houlsby et al. (2019), BitFit Zaken et al. (2022), Prefix Tuning Li and Liang (2021), Prompt Tuning Lester et al. (2021), and Low-Rank Adaptation (LoRA) Hu et al. (2021a). Among them, LoRA is widely adopted for its efficiency in approximating weight updates via low-rank matrices. Extensions such as ReLoRA Lialin et al. (2023) and RankAdapter Zhou et al. (2024) improve memory use and adapt ranks dynamically, though they lack theoretical guarantees. AdaZeta Yang et al. (2024) introduces zeroth-order optimization with convergence guarantees, while others Gao et al. (2024); Rajabzadeh et al. (2024); Valipour et al. (2022) explore adaptive ranks without formal proofs. LoRA has been integrated with Mixture-of-Experts models Li et al. (2024a); Wu et al. (2024a), as in AdaMoLE Liu and Luo (2024), to enable dynamic expert selection, and also with Neural Architecture Search for LLM compression Muñoz et al. (2025). These approaches primarily target weight updates, overlooking direct interventions in hidden representations, which are discussed next.
@@ A.2 Representation Fine-Tuning (ReFT)
ReFT shifts fine-tuning from model weights to hidden representations, leveraging their semantic structure for efficient adaptation Wu et al. (2024b). Inspired by activation steering and representation engineering Avitan et al. (2024); Li et al. (2024b); Liu et al. (2023); Singh et al. (2024), ReFT enables task-specific control through fixed or learned interventions without updating the full model. Notably, Inference-Time Intervention (ITI) Li et al. (2024b) improves LLM truthfulness by modifying activations, while representation engineering Zou et al. (2023) combines representation reading and control for interpretable model behavior. Minimally Modified Counterfactuals (MMC) Singh et al. (2024) unify erasure and steering to reduce bias, and can be mapped to natural language edits Avitan et al. (2024), enhancing interpretability. These findings support direct representation manipulation as a lightweight and effective alternative to weight-based PEFT methods like LoRA.
@@ A.3 Federated Fine-Tuning
Federated Learning (FL) McMahan et al. (2017) poses challenges for fine-tuning LLMs, including data heterogeneity, communication constraints, and model diversity. PEFT methods have emerged to address these issues efficiently Sun et al. (2022); Chen et al. (2022); Zhang et al. (2023). LoRA-based approaches such as FedLoRA Yi et al. (2023), Hyper-FloRA Lu et al. (2024), and Efficient FL Adapter Cai et al. (2023) offer modular and personalized adaptation across clients. Recent advances further incorporate privacy (FFA-LoRA Sun et al. (2024)), heterogeneous adaptation (FloRA Wang et al. (2024)), instruction tuning (FedIT Zhang et al. (2024)), and expert routing (DualFed Long et al. (2024), Sparse-FedMoE Tran et al. (2025)). In contrast, our proposed FedReFT shifts from weight updates to direct representation-level tuning via sparse intervention layers and introduces an All-But-Me (ABM) aggregation strategy to preserve semantic alignment while enabling robust knowledge sharing across non-IID clients.
@@ A.4 Aggregation Methods in FL
To address the inherent heterogeneity and robustness challenges in federated learning, median-based aggregation strategies have been extensively studied as alternatives to simple averaging. Unlike the arithmetic mean, the geometric and coordinate-wise medians are significantly more resilient to outliers and adversarial updates, making them suitable for secure and personalized FL scenarios. For instance, coordinate-wise median aggregation has been proposed to defend against Byzantine clients in distributed optimization Blanchard et al. (2017). This was extended with geometric median-based gradient descent to improve statistical guarantees across diverse loss landscapes Yin et al. (2018). Further work demonstrated that coordinate-wise median and trimmed-mean-based methods achieve order-optimal convergence not only for strongly convex losses but also under non-strongly convex and even non-convex population losses Chen et al. (2017). Additionally, a one-round median-based algorithm was shown to maintain statistical optimality under quadratic convexity, offering a communication-efficient solution Chen et al. (2017). RFA Pillutla et al. (2022) maintains privacy and demonstrates improved robustness over standard averaging techniques, particularly in environments with high levels of data corruption. FedEAT Pang et al. (2025) integrates adversarial training in the embedding space with geometric median-based aggregation to enhance robustness while preserving performance. This work demonstrates that LoRA-based FL systems can effectively leverage geometric median aggregation. Inspired by these findings, we adopt geometric median aggregation in our FL framework to aggregate the All-But-Me (ABM) intervention parameter, weight $\mathbf{W}$, rotation $\mathbf{R}$, and bias $\mathbf{b}$. This provides stability across diverse client behaviors and loss geometries, improving personalization performance under data and objective heterogeneity.
@@ B.1 Hyperparameter Search Space for Commonsense and Arithmetic Reasoning
Following the ReFT framework Wu et al. (2024b), we construct a development set using the GSM8K dataset and consider only the last 300 samples. We trained the clients using LLaMa 7B model with the remaining training data and determined the best-performing hyperparameters based on the model’s performance on the development set. We further use this hyperparameter in another model directly. We set the maximum input sequence length to 512 tokens during training and tuning, and limit inference to 32 generated tokens. We use the same setup for commonsense reasoning with COMMONSENSE170k dataset. The hyperparameter search space is summarized in Tables LABEL:tab:appendix_gsm8k_hypertune and LABEL:tab:appendix_commonsense_hypertune.
During inference, we use greedy decoding (without sampling) for the commonsense reasoning benchmark, as it is a multi-token classification task. For arithmetic reasoning, we follow the decoding setup from  Hu et al. (2023), using a higher temperature of 0.3. This change helps avoid errors in HuggingFace’s decoding caused by unstable probabilities
Table 7: Narrow down the hyperparameter(HP) search space of LLaMA 7B models with FedReFT on the GSM8K development set, inspired from Wu et al. (2024b). The best-performing settings are underlined. We apply greedy decoding without sampling during hyperparameter tuning.
@@ B.2 Hyperparameter Search Space for GLUE Benchmark
We perform hyperparameter (HP) tuning on RoBERTa-large separately for each GLUE task, selecting the optimal settings based on validation performance using a fixed random seed of 42. Final evaluations are conducted using two additional unseen seeds, {43, 44}, to ensure robustness. Table LABEL:tab:appendix_roberta_hyper depicts this.
Table 9: Hyperparameter(HP) settings of RoBERTa-large models on selected GLUE tasks for FedReFT, inspired from Wu et al. (2024b)
@@
Appendix C Dataset Description
C.1 Commonsense Reasoning

### v2

@@ A.1 Parameter-Efficient Fine-Tuning (PEFT)
Fine-tuning LLMs is resource-intensive due to their large parameter counts. Parameter-efficient fine-tuning (PEFT) methods mitigate this by updating only a small subset of parameters while keeping pre-trained weights frozen li2021prefix; he2021towards; wang2022adamix. Several PEFT approaches have been proposed, Adapter Tuning houlsby2019parameter, BitFit zaken2022bitfit, Prefix Tuning li2021prefix, Prompt Tuning lester2021power, and Low-Rank Adaptation (LoRA) Hu21. Among them, LoRA is widely adopted for its efficiency in approximating weight updates via low-rank matrices. Extensions such as ReLoRA lialin2023relora and RankAdapter zhou2024rankadaptor improve memory use and adapt ranks dynamically, though they lack theoretical guarantees. AdaZeta yang2024adazeta introduces zeroth-order optimization with convergence guarantees, while others gao2024adaptive; rajabzadeh2024qdylora; valipour2022dylora explore adaptive ranks without formal proofs. LoRA has been integrated with Mixture-of-Experts models li2024mixlora; wu2024mixture, as in AdaMoLE liu2024adamole, to enable dynamic expert selection, and also with Neural Architecture Search for LLM compression munoz2025low. These approaches primarily target weight updates, overlooking direct interventions in hidden representations, which are discussed next.
@@ A.2 Representation Fine-Tuning (ReFT)
ReFT shifts fine-tuning from model weights to hidden representations, leveraging their semantic structure for efficient adaptation Wu2024. Inspired by activation steering and representation engineering avitan2024changed; li2024inference; liu2023context; singh2024mimic, ReFT enables task-specific control through fixed or learned interventions without updating the full model. Notably, Inference-Time Intervention (ITI) li2024inference improves LLM truthfulness by modifying activations, while representation engineering zou2023representation combines representation reading and control for interpretable model behavior. Minimally Modified Counterfactuals (MMC) singh2024mimic unify erasure and steering to reduce bias, and can be mapped to natural language edits avitan2024changed, enhancing interpretability. These findings support direct representation manipulation as a lightweight and effective alternative to weight-based PEFT methods like LoRA.
@@ A.3 Federated Fine-Tuning
Federated Learning (FL) mcmahan2017communication poses challenges for fine-tuning LLMs, including data heterogeneity, communication constraints, and model diversity. PEFT methods have emerged to address these issues efficiently sun2022exploring; chen2022fedtune; zhang2023fedpetuning. LoRA-based approaches such as FedLoRA yi2023fedlora, Hyper-FloRA lu2024hyperflora, and Efficient FL Adapter cai2023efficient offer modular and personalized adaptation across clients. Recent advances further incorporate privacy (FFA-LoRA sun2024improving), heterogeneous adaptation (FloRA wang2024flora), instruction tuning (FedIT zhang2024towards), residual learning in FRLoRA  yanfederated which tackles client drift by directly adding residual low-rank weight products to the global model parameters in each round, and heterogeneous resources in FlexLoRA  bai2024federated which mitigates the bucket effect by leveraging SVD to aggregate and redistribute LoRA weights with varying ranks among clients, and model compression in FedBiOT wu2024fedbiot which enables LLM fine-tuning without requiring clients to access the full model by using a bi-level optimization scheme to align a compressed emulator with a lightweight adapter, and adaptive sketching in FSLoRA  fang2025federated which reduces communication and computation overhead by allowing clients to selectively update submatrices of the global LoRA modules based on sketching ratios, and expert routing (DualFed long2024dual, Sparse-FedMoE tran2025revisiting). In contrast, our proposed FedReFT shifts from weight updates to direct representation-level tuning via sparse intervention layers and introduces an All-But-Me (ABM) aggregation strategy to preserve semantic alignment while enabling robust knowledge sharing across non-IID clients.
@@ A.4 Aggregation Methods in FL
To address the inherent heterogeneity and robustness challenges in federated learning, median-based aggregation strategies have been extensively studied as alternatives to simple averaging. Unlike the arithmetic mean, the geometric and coordinate-wise medians are significantly more resilient to outliers and adversarial updates, making them suitable for secure and personalized FL scenarios. For instance, coordinate-wise median aggregation has been proposed to defend against Byzantine clients in distributed optimization blanchard2017machine. This was extended with geometric median-based gradient descent to improve statistical guarantees across diverse loss landscapes yin2018byzantine. Further work demonstrated that coordinate-wise median and trimmed-mean-based methods achieve order-optimal convergence not only for strongly convex losses but also under non-strongly convex and even non-convex population losses chen2017distributed. Additionally, a one-round median-based algorithm was shown to maintain statistical optimality under quadratic convexity, offering a communication-efficient solution chen2017distributed. RFA pillutla2022robust maintains privacy and demonstrates improved robustness over standard averaging techniques, particularly in environments with high levels of data corruption. FedEAT pang2025fedeat integrates adversarial training in the embedding space with geometric median-based aggregation to enhance robustness while preserving performance. This work demonstrates that LoRA-based FL systems can effectively leverage geometric median aggregation. Inspired by these findings, we adopt geometric median aggregation in our FL framework to aggregate the All-But-Me (ABM) intervention parameter, weight $\mathbf{W}$, rotation $\mathbf{R}$, and bias $\mathbf{b}$. This provides stability across diverse client behaviors and loss geometries, improving personalization performance under data and objective heterogeneity.
@@ B.1 Hyperparameter Search Space for Commonsense and Arithmetic Reasoning
Following the ReFT framework Wu2024, we construct a development set using the GSM8K dataset and consider only the last 300 samples. We trained the clients using LLaMA 7B model with the remaining training data and determined the best-performing hyperparameters based on the model’s performance on the development set. We further use this hyperparameter in another model directly. We set the maximum input sequence length to 512 tokens during training and tuning, and limit inference to 32 generated tokens. We use the same setup for commonsense reasoning with Commonsense170K dataset. The hyperparameter search space is summarized in Tables LABEL:tab:appendix_gsm8k_hypertune and LABEL:tab:appendix_commonsense_hypertune.
During inference, we use greedy decoding (without sampling) for the commonsense reasoning benchmark, as it is a multi-token classification task. For arithmetic reasoning, we follow the decoding setup from  hu2023llm, using a higher temperature of 0.3. This change helps avoid errors in HuggingFace’s decoding caused by unstable probabilities
Table 7: Narrow down the hyperparameter(HP) search space of LLaMA 7B models with FedReFT on the GSM8K development set, inspired from Wu2024. The best-performing settings are underlined. We apply greedy decoding without sampling during hyperparameter tuning.
@@ B.2 Hyperparameter Search Space for GLUE Benchmark
We perform hyperparameter tuning following common practice for PEFT methods hu2023llm in FL on RoBERTa separately for each GLUE task in Table LABEL:tab:appendix_roberta_hyper, selecting the optimal settings based on validation performance using a fixed random seed of 42. Final evaluations are conducted using two additional unseen seeds, {43, 44}, to ensure robustness.
Table 9: Hyperparameter(HP) settings of RoBERTa-large models on selected GLUE tasks for FedReFT, inspired from Wu2024
@@
Appendix C Theoretical Foundation: Geometric Median via Weiszfeld’s Algorithm
This optimization is non-smooth and convex, and generally lacks a closed-form solution. However, Weiszfeld’s algorithm weiszfeld1937point provides an efficient iterative method to approximate $\mathbf{y}^{*}$. We now derive and justify this algorithm via the Majorization-Minimization (MM) framework.
We define the cost function to be minimized: This function is convex but non-differentiable at points where $\mathbf{y}=\mathbf{x}_{i}$. Weiszfeld’s algorithm avoids such points during updates by construction.
Solving the above yields the Weiszfeld update rule:
The update is only valid when $\mathbf{y}^{(k)}\neq\mathbf{x}_{i}$ for all $i$, a condition that can be enforced by initialization and step-size dampening if needed. From MM theory lange2000optimization, each iteration satisfies:
$f(\mathbf{y}^{(k+1)})$ $\leq Q(\mathbf{y}^{(k+1)}|\mathbf{y}^{(k)})$ (15)
$\leq Q(\mathbf{y}^{(k)}|\mathbf{y}^{(k)})=f(\mathbf{y}^{(k)}),$
C.1 All-But-Me (ABM) Aggregation Strategy
We compute this using Weiszfeld’s algorithm for each parameter type independently, ensuring robustness to outlier clients and misaligned updates. This enables stable and personalized aggregation without sacrificing task-specific semantics. Weiszfeld’s algorithm provides a theoretically grounded and computationally efficient way to compute the geometric median, making it ideal for ABM aggregation in heterogeneous FL. By leveraging this algorithm in FedReFT, we ensure robustness in aggregation and improve both convergence and personalization in non-i.i.d. federated environments.
C.2 Geometric Median over mean
The Geometric Median used in All-but-Me (ABM) aggregation is computationally more expensive, with a time complexity of $\mathbf{O(T\cdot d)}$ due to its iterative Weiszfeld optimization and the need to accumulate all $N$ client updates. In this paper, the dimension is only $d=20$, which makes the overhead relatively small. In contrast, the arithmetic mean used in FedAvg has a lower complexity of $\mathbf{O(d)}$. However, the slightly higher cost of the Geometric Median is justified by its stronger robustness: by minimizing the $\mathbf{L_{1}}$ error, it becomes significantly more resilient to client heterogeneity and outlier updates. This leads to consistently higher accuracy, with clear performance improvements observed over FedAvg on the Commonsense task. While outlier clien updates can render the aggregated FedAvg model unstable or even unusable, the Geometric Median ensures a stable and reliable update, producing a more accurate final model.
For a heterogeneous task setting of a commonsense task using 8 clients, the Geometric Median provides superior outlier detection in Figure 3. The Distance from GeoMed plot visually proves Client 7 is an extreme outlier, clearly separating it from the benign majority’s distance. This maximum separation occurs because GeoMed anchors near the majority by $\mathbf{L}_{1}$ linear error minimization, resisting the outlier’s pull. Conversely, the Mean plot shows the Mean center is contaminated, minimizing the outlier’s distance because the Mean minimizes $\mathbf{L}_{2}$ squared error in figure 4. This distortion makes the aggregated model unusable, but GeoMed’s robust calculation ensures a stable, accurate final update.
Figure 3: Study_Distance_from_GeoMed_ShapeColor
Figure 4: Study_Distance_from_Mean_ShapeColor
Appendix D Ablation Study
D.1 Intervention Parameter Sharing Strategy
To reduce communication overhead while maintaining personalization, we explore three strategies for sharing local intervention parameters with the server. These strategies represent different trade-offs between expressiveness and communication efficiency:
  • Full Intervention Sharing: $\mathit{W}\in\mathbb{R}^{r\times d}$, $\mathit{R}\in\mathbb{R}^{r\times d}$, $\mathit{b}\in\mathbb{R}^{r}$} This strategy shares the complete set of intervention parameters, capturing client-specific compression ($\mathit{W}$), transformation ($\mathit{R}$), and translation ($\mathit{b}$). It enables the most accurate reconstruction of local updates and yields the best global performance, especially under high heterogeneity.
  • No Bias Sharing: $\mathit{W}\in\mathbb{R}^{r\times d}$, $\mathit{R}\in\mathbb{R}^{r\times d}$} This variant omits the bias term $\mathit{b}$ but retains the directional transformation via $\mathit{W}$ and $\mathit{R}$. While it allows the server to align low-rank subspace transformations across clients, it lacks the ability to model per-dimension translation shifts, which can hinder fine-grained personalization.
  • No $\mathit{W}$ Sharing: $\mathit{R}\in\mathbb{R}^{r\times d}$, $\mathit{b}\in\mathbb{R}^{r}$}. This configuration excludes $\mathit{W}$, giving the server access only to the reconstruction and shift parameters. Without knowledge of how the local signals were encoded, the server’s ability to interpret or align updates is severely limited.
The $\{\mathit{W},\mathit{R},\mathit{b}\}$ strategy provides the highest fidelity for aggregation, $\{\mathit{W},\mathit{R}\}$ offers a balanced compromise, and $\{\mathit{R},\mathit{b}\}$ prioritizes communication efficiency at the cost of semantic alignment and global performance.
Table 10: Performance vs. parameter efficiency for different LoReFT sharing strategies (Uplink) for $C$ clients on commonsense reasoning task following the second experiment design.
[…]
GLUE W,R,b 0.01384 94.91
ROBERTa W,R 0.01383 64.03
R,b 0.00693 74.12
Arithmetic W,R,b 0.03114 33.31
LLaMA-2 7B W,R 0.03114 26.13
R,b 0.01557 25.77
Commonsense W,R,b 0.03114 76.55
LLaMA-2 7B W,R 0.03114 70.63
R,b 0.01557 68.02
D.2 Partial Client Participation (PCP)
The ability of a federated learning system to maintain performance when only a fraction of clients participate in each round, known as Partial Client Participation (PCP), is crucial for real-world deployment. This robustness to PCP demonstrates the system’s ability to handle practical constraints like communication bandwidth limitations, client availability, and device battery life. To evaluate partial client participation (PCP), the experiment’s results in Table LABEL:tbl:glue_task_pcp use a total of $C=5$ clients. When all five clients participated in every global round, FedReFT obtained an average accuracy of $91.45\%$. Under PCP, where only $C=3$ clients were randomly selected per round, FedReFT achieved an average accuracy of $90.62\%$. The small difference between the two settings indicates that FedReFT maintains stable performance even with reduced client participation.
Table 11: Partial client participation (PCP) and all clients in FL settings for GLUE Task for FedReFT
Method MNLI-m SST-2 QNLI QQP Avg $\uparrow$
PCP 89.17 92.83 93.81 86.65 90.62
All 89.61 95.25 94.02 86.93 91.45
D.3 Integrating Test-Time Computing (TTC) with Adaptive Mixing in Local Model Updates
In FedReFT, local interventions encode client-specific semantics, while ABM aggregation conveys shared global knowledge. However, naively averaging representation-level updates can disrupt semantic consistency across clients, leading to misaligned feature spaces. Furthermore, uniform aggregation disregards client heterogeneity, reducing adaptability in diverse edge environments.
To address these challenges, we evaluate the effectiveness of FedReFT under two mixing strategies: (i) adaptive mixing via Test-Time Computing (TTC), where local and global interventions are dynamically balanced, and (ii) balanced mixing, where a fixed coefficient $\alpha$ governs the trade-off between local ($\alpha$) and global ($1-\alpha$) contributions. While TTC provides a dynamic, task-adaptive mechanism expected to outperform balanced mixing ($\alpha=0.5$), we also include results for the fixed $\alpha=0.5$ setting to illustrate this trade-off. Table LABEL:tbl:baseline_first_commmonsense_ablation, LABEL:tbl:glue_task_ablation, 14, and LABEL:tbl:math_setup_3_baseline_ablation show that the performance of TTC-based adaptive mixing is over the balanced mixing.
In equation 9, the Test-Time Computing (TTC) optimization uses three small regularization weights, $\lambda_{1},\lambda_{2},\lambda_{3}$, all set to 0.001. This low value means the primary goal is minimizing the test loss $\mathcal{L}_{\text{test}}$, not strict following of the regularization rules. $\lambda_{1}$ Entropy and $\lambda_{3}$ Diversity gently stop the mixing coefficient $\alpha$ from completely favoring local (1) or global (0) updates. $\lambda_{2}$ Consistency mildly allows $\alpha$ to change across different intervention layers, which helps with adaptive personalization. Because the $\lambda$ values are small, the resulting mean $\alpha\approx 0.57$ strongly favors local updates based on the client’s specific data, but not the fully Local-Only assumption of $\mathbf{\alpha=}$1.0. If these $\lambda$ values were increased significantly, the regularization would become more important than the test loss. This would push $\alpha$ closer to $0.5$ for balanced mixing across all layers, reducing the model’s ability to personalize.
Table 12: Federated fine-tuning performance of LLaMA-3.2 3B across five commonsense reasoning tasks with the Mixed-Task (MT) setup, where clients train on heterogeneous task mixtures. We report the effectiveness of FedReFT under two mixing strategies: (i) adaptive mixing using Test-Time Computing (TTC), where local and global interventions are dynamically balanced, and (ii) balanced mixing, where a coefficient $\alpha$ controls the trade-off ($\alpha$ for local and $1-\alpha$ for global). While TTC is intended to provide a dynamic, task-adaptive solution expected to surpass balanced mixing ($\alpha=0.5$), we also include results for $\alpha=0.5$ to illustrate the trade-off. R = Rank, TP(M) = Trainable Parameters in Millions, and TP Effi. (R8) = parameter efficiency of FedReFT with rank 8 compared to baselines. Baseline numbers are taken from singhal2025fed.
[…]
16 5.5 $0.5\times$ 63.42 81.61 73.64 91.23 71.35 76.05
32 11.01 $0.25\times$ 64.53 81.34 73.39 91.51 71.32 76.22
FedReFT (tie $\phi$) 4 0.688 0.0214 49.94 81.23 72.72 89.84 68.43 72.43
8 1.38 0.0428 57.15 81.22 72.77 90.56 68.50 74.04
Table 13: Performance comparison across GLUE tasks on RoBERTa with $C=3$. We report results under two mixing strategies: (i) adaptive mixing with Test-Time Computing (TTC), which dynamically balances local and global interventions at inference, and (ii) balanced mixing, where $\alpha$ denotes the proportion of local contribution and $1-\alpha$ the global contribution. The balanced $\alpha=0.5$ cases illustrate trade-offs between local personalization and global knowledge, while TTC is intended as a dynamic alternative expected to generalize better across heterogeneous conditions. TP denotes the number of trainable parameters (in millions), and TP Effi. measures the parameter efficiency of FedReFT (rank 1) relative to all LoRA-based baselines guo2024selective with rank 8.
[…]
Table 14: Performance comparison across arithmetic reasoning tasks with the Distinct Task (DT) and Mixed Task (MT) setup using different model sizes. We report results under two mixing strategies: (i) adaptive mixing with Test-Time Computing (TTC), which dynamically balances local and global interventions at inference, and (ii) balanced mixing, where $\alpha=0.5$ denotes the proportion of local contribution and $1-\alpha$ the global contribution.
Distinct Task (DT) Mixed Task (MT)
FedReFT ($\alpha=0.5$) AQuA SVAMP MAWPS Avg $\uparrow$ AQuA SVAMP MAWPS Avg $\uparrow$
LLaMA 7B 25.59 25.47 49.80 33.62 22.83 14.33 27.10 21.42
LLaMA-2 7B 29.53 32.45 57.30 39.76 21.65 20.39 31.50 24.51
LLaMA-3 8B 34.64 48.98 73.60 52.41 31.89 48.90 70.04 50.48
FedReFT (TTC) Distinct Task (DT) Mixed Task (MT)
LLaMA 7B 26.12 26.71 49.94 34.26 20.86 15.60 28.22 22.23
LLaMA-2 7B 30.96 33.31 59.51 41.93 22.05 23.50 32.74 26.09
LLaMA-3 8B 35.36 49.41 75.25 53.34 33.45 51.28 73.51 52.75
D.4 Comparison of Aggregation Methods on Different Tasks
We use training data from the COMMONSENSE170K dataset, split among three clients, and evaluate the models using the SIQA task. Figure 5 shows that Geometric Median ABM aggregation outperforms all other approaches. Similarly, we split the MATH10K dataset among three clients, train each client for only five local epochs, and evaluate the results using the GSM8K evaluation set. Additionally, we use the QNLI GLUE dataset for the natural language understanding (NLU) task. Why small gains matter. We emphasize that even modest accuracy improvements are meaningful in the federated learning setting, particularly under highly heterogeneous task distributions. As shown in Figure 5 the arithmetic mean has time complexity $O(d)$, while the geometric median (Weiszfeld’s algorithm) has $O(T\cdot d)$, with $T$ as iterations and $d$ as parameters. Both methods have memory complexity $O(d)$. Despite the higher computational cost, the geometric median provides greater robustness to heterogeneity and consistently yields higher accuracy in FL settings.
[…]
Figure 5: Comparison of aggregation strategies across tasks. Results for FedAvg, Mean-ABM, and Geometric Median-ABM on Commonsense Reasoning, Arithmetic Reasoning, and GLUE show that Geometric Median-ABM consistently outperforms others, demonstrating greater robustness in heterogeneous federated settings.
These improvements hold consistently across three diverse tasks. We chose Geometric Median-ABM not only for its peak accuracy but also for its robustness to outliers and task drift. The benefits are especially pronounced in arithmetic reasoning, where client diversity is highest. While geometric median incurs additional server-side computation, it is applied once per round on low-dimensional sparse parameters, making the overhead negligible. Client-side operations remain lightweight.
$f(\mathbf{y})=\sum_{i=1}^{n}\|\mathbf{y}-\mathbf{x}_{i}\|_{2}.$ (17)
Appendix E Communication Efficiency
E.1 Intervention Parameter sharing Across Tokens in Same Layer
In this section, we also conducted some additional experiments to show the robustness of FedReFT in different setups. We experiment with whether to share (tie) the intervention parameters $\phi$ across different input positions within the same layer. Given the positions $P=\{1,\ldots,p\}\cup\{n-s+1,\ldots,n\}$, we define the untied and tied variants Wu2024:
while FedReFT (tie $\phi$) offers a compelling trade-off between performance and efficiency. These results highlight the scalability and efficiency of our representation-tuning approach. Appendix Table 12 and 14 depicts these.
Table 16: We vary LLaMA model sizes with $C=3$ clients following the Distinct Task (DT) design for the commonsense reasoning task, alongside a centralized LoReFT baseline. As model capacity increases, we observe notable performance gains, with the largest model approaching the accuracy of the centralized setting.
Method BoolQ PIQA HellaS. Avg $\uparrow$
Tiny LLaMA 1B 63.92 50.36 47.21 53.83
LLaMA 7B 66.64 78.34 67.92 70.97
LLaMA-2 7B 69.71 75.45 78.26 74.47
LLaMA-3.2 3B 65.93 78.37 82.36 75.55
E.2 Additional Experimental Validation
We experiment with different LLaMA model sizes across $C=3$ clients following the Distinct Task (Dt) framework for the commonsense reasoning task in Table 16, along with a centralized LoReFT baseline for comparison. As the model size grows, we observe steady performance improvements, with the largest variant achieving results close to the centralized model. The first four experiments are conducted under the standalone (centralized) setup, and the following four are performed in the federated learning (FL) environment.
[…]
LLaMA-1.1B 1100.05 0.72 0.0655
LLaMA 7B 6,738.42 2.10 0.0311
LLaMA-2 7B 6,738.42 2.10 0.0311
LLaMA-3 8B 8,030.27 2.10 0.0261
LLaMA-2-13B 13,015.86 6.55 0.0503
RoBERTa Large 355.36 0.050 0.0138
Appendix F Dataset Description
F.1 Commonsense Reasoning